OpenStudy (anonymous):

Medal & Fan!!!! Prove: tan^2(θ)•cos^2(θ) + cos^2(θ) = 1. You must show all work

OpenStudy (midhun.madhu1987):

= cos^2(θ (tan^2(θ) + 1) = cos^2 x sec^θ =1

OpenStudy (anonymous):

could you please describe how u got that?? @midhun.madhu1987

OpenStudy (midhun.madhu1987):

We have an Identity in trigonometry: sec^θ = tan^2(θ) + 1